40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 30 18:44:11 2021, Anchal, Summary, LSC, Tried fixing ETMY QPD 
    Reply  Thu Jul 1 10:34:51 2021, Paco, Anchal, Summary, LSC, ETMY QPD fixed 
Message ID: 16233     Entry time: Thu Jul 1 10:34:51 2021     In reply to: 16232
Author: Paco, Anchal 
Type: Summary 
Category: LSC 
Subject: ETMY QPD fixed 

Paco worked on alignign the beam splitter to get light on the ETMY QPD and was successful in centering it without any other changes in the settings.

ELOG V3.1.3-